OpenStudy (aonz):

The expression 3x^2 + 2xy − 8y^2 − 8x + 14y −3 can be treated as a function in x with y as an arbitrary constant or as a function in y with x as an arbitrary constant. Show that in either case the expression is indefinite. Factor the expression.

OpenStudy (anonymous):

@AonZ By indefinite is the question saying that in either case the function approaches infinity as x or y approach infinity?

OpenStudy (aonz):

i dont think so. Indefinite means you that \[\Delta \ge0\]

OpenStudy (unklerhaukus):

\[\Delta=b^2-4ac\]?

OpenStudy (unklerhaukus):

if y is a constant \[(3)x^2 + (2y-8)x +( 14y- 8y^2 −3)\] \[(a)x^2+(b)x+(c)\]

OpenStudy (aonz):

yea i done that then i got 25y^2 - 34y + 7 when i used the discriminant

OpenStudy (anonymous):

With the same approach, if 'x' is constant then we get:\[\bf (-8)y^2+(14+2x)y+(3x^2-8x-3)\]\[\bf -8y^2+(d)y+(e)\]

OpenStudy (anonymous):

@AonZ So \(\bf \Delta\) does represent the discriminant?

OpenStudy (aonz):

yes

OpenStudy (anonymous):

Ok so in each case, we'll have to consider the discriminant separately. I'll first consider the case in which 'x' is considered a constant. In such a case, the discriminant is:\[\bf b^2-4ac=(14+2x)^2-4(-8)(3x^2-8x-3)\]\[\bf =196+56x+4x^2+96x^2-256x-96=100x^2-200x+100\]\[\bf =100(x^2-2x+1)=100(x-1)^2= \Delta \ge 0\]So we have proved the discriminant is always greater than 0 when we consider 'x' as a constant. Now you must do the same when we consider 'y' a constant. @AonZ

OpenStudy (aonz):

oh true... then after we proved it is always greater of less than 0 we factorise the expression?

OpenStudy (anonymous):

I believe that for the factoring part they are asking you to factor the original equation that is given..

OpenStudy (aonz):

ok, that shouldnt be too hard. Thanks for your help!!!

OpenStudy (anonymous):

yw